The speed of wave on a string with a fixed end and a free end is 300 m/s. The frequency of a wave is 250Hz. Find the length of the string necessary to produce a standing wave with first harmonic.
1
Expert's answer
2020-05-25T10:57:54-0400

The wavelength of a standing wave

"\lambda=\frac{v}{f}=\frac{300\:\rm m/s}{250\:\rm Hz}=1.2\:\rm m"

The length of the string

"L=\frac{1}{4}\lambda=\frac{1}{4}\times 1.2\:\rm m=0.3\: m"